One number is 8 less than another. If 5 times the smaller number minus 2 times the larger number is 14, find the two numbers.
Answer by fractalier(2101)

Call the larger number x.
The smaller is then x - 8.
From the English in the problem, we convert to an algebraic equation...
5(x - 8) - 2x = 14
5x - 40 - 2x = 14
3x - 40 = 14
3x = 54
x = 18
x - 8 = 10
and we're done...